The perimeter of the rectangle is 190 inches. The length exceeeds the width by 53 inches. Find the length and the width.
Can someone please help with this one please? The perimeter is the 190 diveded by 4? Is that the way to start or not?

Answer by jojo14344(1513)

Dividing the perimeter by 4? You have percentage of being right but not really the way to start it.
Okay,let's get the formula for perimeter first (P) = 2(L+W), right? This is equation 1.
Knowing the length(L) exceeds the width(W) by 53, so L=W+53, right? and mark this equation 2. Substitute this value to the 1st eq'n knowing the perimeter:
190=2(W+53+W)=2(2W+53)=4W+106
190-106=4W
84/4=W
21inches=Width
Go back eq'n 2: L=21+53
L=74 inches.
Simply to check if in doubt,
P=2(L+W)
190=2(21+74)=2(95)
190=190 cool!
thank you,
Jojo